The problem is to solve the quadratic equation using the Indian method of multiplying each term in the equation by four times the coefficient of the
x sqrt 2 term.
The equation is: x^ - 2x - 13 = 0
I am totally stumped. Please help.
Ellen

Answer by oscargut(2103)

x^2 - 2x - 13 = 0
x^2 - 2x =13
4x^2-8x=52
4x^2-8x+4=56
(2x-2)^2=56
2x-2=sqrt(56) or 2x-2=-sqrt(56)
x=(sqrt(56)+2)/2 or x=(-sqrt(56)+2)/2